OpenStudy (anonymous):

area between 2 curves. please help

OpenStudy (anonymous):

I have to find the area between \[x=y^2+1\]\[and\]\[x=9-y^2\]

OpenStudy (anonymous):

this is the picture I got

OpenStudy (anonymous):

but I have 1 question

OpenStudy (anonymous):

will the area below the x-axis cancel the area above the x-axis ?

OpenStudy (zale101):

Graph them both and see what is the area between the curves that overlaps.

OpenStudy (anonymous):

so, the area below the x-axis has to be canceling the area above the x-axis in any case where I am finding area between 2 curves?

OpenStudy (anonymous):

if so the answer is 0, and if not, then...

OpenStudy (anonymous):

if not, I will re-write the functions in terms of function y(x), integrate the first from 1 to 5, and the second from 5 to 9, and add the areas (including only positive root) and then mutiply the sum times 2

OpenStudy (anonymous):

yes, I got the desmos picture

OpenStudy (anonymous):

\[y=\sqrt{x-1}\]\[y=\sqrt{9-x}\]\[A(x)=2\left(\int\limits_{1}^{5} (\sqrt{x-1})dx~+~\int\limits_{5}^{9}(\sqrt{9-x})~dx\right)\]

OpenStudy (anonymous):

this is if the area under the x-axis does NOT cancel with the area above the x-axis.

OpenStudy (anonymous):

and the part I don't know, is whether the area below and above x-axis cancel each other or not,

OpenStudy (zale101):

What value did you get after solving the definite integral?

OpenStudy (anonymous):

I am too lazy now, I will put it into wolfram

OpenStudy (zale101):

I will see if the value is zero or not.

OpenStudy (anonymous):

well, I did the area above the x-axis times two

OpenStudy (anonymous):

so that is positive.

OpenStudy (anonymous):

and again, this is where my problem is. Is it correct to just multiply the area bounded by the two curves and the x-axis times 2, or do area below and above the x-axis cancel. do you know ?

OpenStudy (zale101):

How come you multiplied the integrals by 2?

OpenStudy (zale101):

lol sorry, i'm just a little lost here. Just covered this topic yesterday xD

OpenStudy (anonymous):

the integrals inside the parenthesis are the area between the 2 curves that is above the x-axis. Correct ?

OpenStudy (anonymous):

then, there is the exact mirror/reflection of this shape/area below the x-axis.

OpenStudy (anonymous):

or lost?

OpenStudy (anonymous):

https://www.desmos.com/calculator/zsteb4lgbv this is I re-wrote the functions in terms of y=f(x) BUT without including the +- by the square roots, which therefore gives me just the area above the x-axis.

OpenStudy (anonymous):

and then, after finding all the area between the curves which is above the x-axis, I multiply the result times 2. Why do multiply times 2, because there are 2 exactly same regions. 1) below x-axis 2) above x-axis

OpenStudy (anonymous):

this is all, IFF areas above and below x-axis don't cancel each other. I know how to calculate, but now what to calculate... what a pain -:(

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

tagged:)

OpenStudy (zale101):

 X=y^2+1 x=9-x^2 \[dA=(X-x)dy\] \(dA=(y^2+1-9-x^2)dy\) \[\int\limits_{-2}^{2}(y^2+1-9-y^2)dy\] I chose to do with the respect to y, you can also do with the respect to x